1. Two electrons are separated by a distance of 5.6×10^-11 m in air.


a. Find the magnitude of the electrical force one electron exerts on the other. Is this force attractive or repulsive?

b. Find the magnitude of the gravitational force one electron exerts on the other. Is this force attractive or repulsive?


1
Expert's answer
2022-03-29T12:27:37-0400

(a)

"F=k\frac{e^2}{r^2}\\
=9*10^9*\frac{(1.6*10^{-19})^2}{(5.6*10^{-11})^2}=7.3*10^{-8}\:\rm N"


repulsive

(b)

"F=G\frac{m^2}{r^2}\\
=6.67*10^{-11}*\frac{(9.1*10^{-31})^2}{(5.6*10^{-11})^2}=1.3*10^{-50}\:\rm N"

attractive